# gen2seg: Generative Models Enable Generalizable Instance Segmentation

### v1

@@ 4.4 Zero-Shot Promptable Instance Segmentation
Figure 7: Delineating the edges of the objects in the scene is a fundamentally ambiguous task. While our models’ outputs do not exactly match the ground truth (neither do SAM’s), they represent one interpretation of the "objects" in the scene. Our model tends to include certain objects, such as the clouds or grass, in the background. This emerges without supervision and may be an inherent bias from generative pretraining.
4.5 Zero-Shot Edge Detection
We evaluate our segmentation features on the task of edge detection with the BSDS500 dataset [41]. It is important to note that we are not simply trying to find all edges in the image, but only object boundaries. Following [29], we simply apply a Sobel filter on the predicted features followed by nonmax suppression to thin the edges. We use the edge detection described above on the original image as a “weak” baseline and on the output of SAM’s AutoMaskGenerator as a “strong” baseline.
Results. Our model accurately delineates the edges of primary objects in each image despite never seeing a mask of the object type. As shown in Figure 6, our Stable Diffusion variant outperforms SAM in the first quarter of the precision-recall curve, and our MAE ViT-H variant matches SAM for the first 15% of recall despite having several orders of magnitude less pre-training. However, as visualized in Figure 7, many labeled regions exist at the interface between object and background (e.g., clouds, plants, or rocks). Our model tends to include these regions in the background. As a result, no edges are detected for some objects in our features, and the precision falls for higher recalls. This behavior is fully emergent, as our model has never seen masks of the overwhelming majority of objects present in the BSDS500 dataset. This hypothesis is further supported by how SD-Limited slightly outperforms SD. Specifically, the SD-Limited has no loss computed on a high amount of the background (because we ignore loss on pixels inside the bounding box of unknown-category objects). As a result, our SD-Limited sometimes activates on objects that SD would consider background.
5 Conclusion
Our findings suggest that learning to synthesize visual reality inherently teaches a detailed understanding of its constituent parts. Our models are able to segment objects and styles unseen in pretraining and finetuning, yet still achieve competitive performance with heavily supervised models. As we continue to scale generative models and diversify their pretraining data, their ability to perceive the visual world will only grow. Learning to leverage these powerful representations for a wide variety of visual tasks has the potential to enable a new frontier in generalizable perception.
A Experimental Models and Datasets
A.1 Additional Details for Models and Baselines
Stable Diffusion v2 (SD): We finetune Stable Diffusion U-Net end-to-end using our loss. To train it in pixel space, we fix the timestep to the highest (999). We replace the input with our image’s VAE embedding instead of the noisy latent. We set the CLIP embedding to the null condition. We freeze the VAE when finetuning. Following [15], this enables deterministic one-step prediction.
Stable Diffusion v2 - Category Limited (SD-Limited) About 30% of masks in the Hypersim dataset have no labeled category. While we have manually verified this to be true, it remains impossible to empirically prove there is no mask of any given category (i.e. dolphin) in our dataset. Thus, we finetune Stable Diffusion v2 as described earlier, but we compute no loss on any pixels inside the bounding box of any mask with the "unknown" category. This enables us to quantitatively show that our model generalizes even in a strict, category limited setting.
DINO + VAE (DINO-B): We investigate whether a strong discriminative model can suffice for segmentation features without necessarily learning to synthesize images. However, we must pair it with a model that knows how to upsample images of all types from a low-dimensional space so that it will be able to generate images of objects unseen in training. To investigate this, we join ImageNet-1K pretrained DINO ViT-B with the Stable Diffusion VAE, and connect them with a single up-conv layer. Additionally, the latent features of images compressed with the Stable Diffusion VAE often look like the image itself [30]. Thus, to succeed, all DINO needs to do is generate these object groupings at a lower resolution.
Segment Anything (SAM): SAM is a large-scale promptable instance segmentation model supervised on over a billion masks from a very large distribution of data. We use SAM as a benchmark to evaluate the extent to which our model generalizes across a wide variety of images. While in some cases we outperform SAM, we do not mean for this to claim that our models are inherently superior to it. We use the publicly available checkpoint and inference library.
To match their pretraining resolutions, SimpleClick, DINO, and MAE are finetuned at 224$\times$224 resolution, rather than the standard ones described in implementation details.
A.2 Evaluation Datasets
PIDRay: Humans effortlessly apply their understanding of object shapes even in scenarios never encountered in nature. For instance, TSA agents can quickly identify dangerous items in X-ray images of luggage. To assess our models’ generalization to this challenging scenario, we employ the PIDRay dataset, which features labeled examples of hazardous objects in baggage. This task is particularly demanding because many dangerous items are small and deliberately concealed within other objects. We evaluate on the "easy" subset of the test set.
B Implementation details
We train our models on a node with four RTX6000 Ada GPUs. We train our models with a batch size of 2, with gradient accumulation for 4 steps, for an effective batch size of 32. We do this intentionally, as described in [28], to mix gradients between images sampled from Hypersim and Virtual Kitti 2. We train our model for 30,000 iterations, which takes about 29 hours for SDv2 and 12 hours for MAE ViT-H. However, our models show no signs of overfitting, and performance would likely benefit from additional iterations, but we didn’t explore this due to timing constraints. We sometimes struggle with memory constraints when finetuning Stable Diffusion, as some images in Hypersim have a very high number of instances (2000+). Thus, we compute the loss on up to 1250 instances at most. We normalize final pixel-level outputs as we observe it improves convergence. We start our learning rate at 6e-5, after 100 steps of warmup. We then decay on a cosine schedule so that we end at $\frac{1}{20}$ of our original learning rate. We train at a resolution of 480$\times$640 for Hypersim and 368$\times$1024 for Virtual Kitti 2 for our Stable Diffusion variants. For ImageNet-pretrained models (MAE, DINO, SimpleClick), we resize Hypersim images to 224$\times$224 and randomly crop a 224$\times$224 region in Virtual Kitti 2 dataset. We set $\lambda_{\mathrm{sep}}\text{ and }\lambda_{\mathrm{mean}}$ to 300 and compute all losses in the range of $[0,255]$ to weight all terms equally. For Stable Diffusion, we finetune only the U-net and freeze the VAE. We set the text condition to the empty string. For all prompting experiments, we fix the threshold to $\frac{3}{255}$ and use a window size of 9 (for joint bilateral smoothing).
C Ablation Study
@@
We present an ablation study in Table 2 to examine the effect of each loss component and the VAE normalization using mIoU at a single center point on the $\text{COCO}_{\text{exc}}$ dataset. First, removing the intra-instance variance loss ($\mathcal{L}_{\mathrm{var}}$) (row 1) causes performance to collapse. This is essential to our method; without it, the model does not produce uniform masks. Eliminating the pixel-level separation loss (row 2) prevents the model from learning sharp object boundaries, causing it to slightly overestimate mask boundaries. Eliminating the mean-level loss (row 3) results in a reduced mIoU as well, but instead affects the model’s ability to discriminate smaller objects. Replacing the smooth $\ell_{1}$ loss with $\ell_{2}$ loss (row 4) and the normalization of VAE outputs (Norm, row 5) results in mIoU of 27.7 and 26.2, respectively. Both smooth $\ell_{1}$ loss and normalization of VAE output helps the model converge to lower loss values earlier. Overall, the complete model that integrates all these components (row 6) achieves the highest mIoU of 31.6, confirming that each element plays a role in obtaining optimal performance.
D List of Object Types in Hypersim and $\text{COCO}_{\text{exc}}$
@@ D.1 Hypersim
This list includes only objects with instance and class annotations. Certain objects which have instance-level annotations but lack class labels, such as teddy bears or potted plants, were placed into the "unknown" object category.
@@ D.2 $\text{COCO}_{\text{exc}}$
We excluded object categories that either labeled in our train set or we have observed to appear. For example, we have excluded "wine glass" or "knife" as these would be placed into Hypersim’s "unknown" category. We have personally verified all of the objects below to not exist in the subset of Hypersim we train on.
@@
Image DINO-B MAE-B MAE-H SD-Limited SD SAM
@@
Figure 8: Qualitative Results on the $\text{COCO}_{\text{exc}}$ [36] dataset. These results are randomly chosen and not cherry-picked.
Image DINO-B MAE-B MAE-H SD-Limited SD SAM
@@
Figure 9: Qualitative Results on the DRAM [11] dataset. These results are randomly chosen and not cherry-picked.
Image DINO-B MAE-B MAE-H SD-Limited SD SAM
@@
Figure 10: Qualitative Results on the EgoHOS [64] dataset. These results are randomly chosen and not cherry-picked.
Image DINO-B MAE-B MAE-H SD-Limited SD SAM
@@
Figure 11: Qualitative Results on the iShape [62] dataset. These results are randomly chosen and not cherry-picked.
Image DINO-B MAE-B MAE-H SD-Limited SD SAM
@@
Figure 12: Qualitative Results on the PIDRay [63] dataset. These results are randomly chosen and not cherry-picked.
DINO-B Refer to caption Refer to caption MAE-B
MAE-H Refer to caption Refer to caption SD-Limited
SD Refer to caption Refer to caption SAM
Figure 13: Qualitative comparison for a challenging in-the-wild image, displayed in high resolution.
DINO-B MAE-B MAE-H
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
SD-Limited SD SAM
Figure 14: Qualitative comparison for a challenging in-the-wild image, displayed in high resolution.
DINO-B Refer to caption Refer to caption MAE-B
MAE-H Refer to caption Refer to caption SD-Limited
SD Refer to caption Refer to caption SAM
Figure 15: Qualitative comparison for a challenging in-the-wild image, displayed in high resolution.
DINO-B Refer to caption Refer to caption MAE-B
MAE-H Refer to caption Refer to caption SD-Limited
SD Refer to caption Refer to caption SAM
Figure 16: Qualitative comparison for a challenging in-the-wild image, displayed in high resolution.
DINO-B Refer to caption Refer to caption MAE-B
MAE-H Refer to caption Refer to caption SD-Limited
SD Refer to caption Refer to caption SAM
Figure 17: Qualitative comparison for a challenging in-the-wild image, displayed in high resolution.
DINO-B MAE-B MAE-H
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
SD-Limited SD SAM
Figure 18: Qualitative comparison for a challenging in-the-wild image, displayed in high resolution.
DINO-B Refer to caption Refer to caption MAE-B
MAE-H Refer to caption Refer to caption SD-Limited
SD Refer to caption Refer to caption SAM
Figure 19: Qualitative comparison for a challenging in-the-wild image, displayed in high resolution.

### v2

@@ Appendix B Additional Results and Experimental Details
Figure 7: Delineating the edges of the objects in the scene is a fundamentally ambiguous task. While our models’ outputs do not exactly match the ground truth (neither do SAM’s), they represent one interpretation of the “objects” in the scene. Our model tends to include certain objects, such as the clouds or grass, in the background. This emerges without supervision and may be an inherent bias from generative pretraining.
@@
We explore if our masks improve with additional prompt points (Figure 6). We average IoU across datasets, excluding $\text{COCO}_{\text{exc}}^{M/S}$ (since few small objects are detected in the first place). Remarkably, our SD model achieves over 82% mIoU relative to SAM at 9 prompt points, despite lacking a learned prompt encoder for multi-point prompts or masks from any of the evaluated categories.
Additionally, as shown in Figure 6, our Stable Diffusion variant outperforms SAM in the first quarter of the precision-recall curve, and our MAE ViT-H variant matches SAM for the first 20% of recall despite having a substantially smaller training resolution when evaluated on edge accuracy on BSDS500. However, as visualized in Figure 7, many labeled regions exist at the interface between object and background (e.g., clouds, plants, or rocks). Our model tends to include these regions in the background. As a result, no edges are detected for some objects in our features, and the precision falls for higher recalls. This behavior is fully emergent, as our model has never seen masks of the overwhelming majority of objects present in the BSDS500 dataset. This hypothesis is further supported by how our models with 5 and 10 classes only perform better. Specifically, they have no loss computed on a high amount of the background (because we ignore loss for pixels inside the bounding box of unknown-category objects). As a result, our 5 and 10 category models sometimes activate on objects that SD or MAE-H would consider background. Thus, we opt to report AP under 20% recall as the lower precision later in the curve is not due to inaccurate edges, but rather a lack of segmented objects in those regions.
@@
We present an ablation study in Table 4 to examine the effect of each loss component and the VAE normalization using mIoU at a single center point on the $\text{COCO}_{\text{exc}}$ dataset. First, removing the intra-instance variance loss ($\mathcal{L}_{\mathrm{var}}$) (row 1) causes performance to collapse. This is essential to our method; without it, the model does not produce uniform masks. Eliminating the pixel-level separation loss (row 2) prevents the model from learning sharp object boundaries, causing it to slightly overestimate mask boundaries. Eliminating the mean-level loss (row 3) results in a reduced mIoU as well, but instead affects the model’s ability to discriminate smaller objects. Replacing the smooth $\ell_{1}$ loss with $\ell_{2}$ loss (row 4) and the normalization of VAE outputs (Norm, row 5) results in mIoU of 27.7 and 26.2, respectively. Both smooth $\ell_{1}$ loss and normalization of VAE output helps the model converge to lower loss values earlier. Overall, the complete model that integrates all these components (row 6) achieves the highest mIoU of 31.6, confirming that each element plays a role in obtaining optimal performance. We also ablate the fixed timestep value for our SD model in Table 5.
In Table 6, we explore our models’ robustness to a variety of image perturbations. This helps show that our model is robust to changes in color and texture, despite never receiving any supervision for this.
Finally, in Table 7, we explore how deep we need to finetune our model to achieve strong results. We start with full finetuning and gradually reduce the layers trained to the last N layers in the model, plus the final projection to pixel space, freezing the rest of the model. We find that finetuning just the decoder leads to very similar performance to full finetuning. Additionally, we find that just one or two layers are enough to achieve reasonable performance. This suggests that the ”core” object grouping information is highly saturated in the decoder, especially the last layers. This makes sense because 1) since the rest of the model is frozen, only the representations that are available at the last N layers will actually have an effect on the output 2) these layers are the ones most responsible for synthesizing images, for which understanding object grouping is frozen.
SD timestep $\text{COCO}_{\text{exc}}^{L}$ $\text{COCO}_{\text{exc}}^{M}$ $\text{COCO}_{\text{exc}}^{S}$ DRAM EgoHOS iShape PIDRay
t=999 (original) 57.6 38.8 8.5 48.2 40.0 51.4 30.9
t=499 58.8 37.5 6.2 48.9 28.6 50.9 31.5
t=1 57.9 38.4 8.5 47.8 31.3 49.2 28.0
Table 5: We ablate the role of the fixed timestep in our results. As shown above, it does not make a major difference on the results as the timestep cross-attention is finetuned in the U-net.
SD MAE-H model
Augmentation $\text{COCO}_{\text{exc}}^{L}$ $\text{COCO}_{\text{exc}}^{M}$ $\text{COCO}_{\text{exc}}^{S}$ $\text{COCO}_{\text{exc}}^{L}$ $\text{COCO}_{\text{exc}}^{M}$ $\text{COCO}_{\text{exc}}^{S}$
Original 57.6 38.8 8.5 50.0 23.2 3.5
Solarize 58.6 35.5 7.7 50.8 21.3 1.1
[…]
Grayscale 56.7 34.0 7.7 45.5 23.0 3.1
Hue +0.3 55.8 34.2 7.9 49.7 21.8 1.4
Hue -0.3 57.0 35.9 8.2 51.0 23.2 3.5
Table 6: We evaluate our models’ zero-shot accuracy robustness to several image color changes and perturbations. As shown below, our models exhibit strong robustness to all with limited drop in mask quality. Particularly, the solarize augmentation results loosely suggest that our model is not grouping based on low-level color/texture and has some sense of high-level grouping.
Last N blocks finetuned $\text{COCO}_{\text{exc}}^{L}$ $\text{COCO}_{\text{exc}}^{M}$ $\text{COCO}_{\text{exc}}^{S}$
MAE-H (original, full finetune) 50.0 23.2 3.5
Decoder only (8 layers) 53.5 20.2 3.0
4 layers 45.5 16.3 2.5
2 layers 40.0 13.0 2.0
1 layer 36.4 10.6 1.4
0 layers (linear probe) 19.6 2.7 0.0
Table 7: We explore how many layers need to be finetuned for our method to work. Surprisingly, full finetuning does not yield performance gains over finetuning the decoder. Additionally, just one or two layers need to be finetuned for reasonable performance.
Appendix C Implementation details
[…]
Appendix D List of Object Types in Hypersim and $\text{COCO}_{\text{exc}}$
@@ D.1 Hypersim
This list includes only objects with instance and class annotations. Certain objects which have instance-level annotations but lack class labels, such as teddy bears or potted plants, were placed into the “unknown” object category.
@@ D.2 $\text{COCO}_{\text{exc}}$
We excluded object categories that either labeled in our train set or we have observed to appear. For example, we have excluded “wine glass” or “knife” as these would be placed into Hypersim’s “unknown” category. We have personally verified all of the objects below to not exist in the subset of Hypersim we train on.
@@
Image DINO-B MAE-B MAE-H SD (10 Classes) SD SAM
@@
Figure 8: Qualitative Results on the $\text{COCO}_{\text{exc}}$ (Lin et al., 2014) dataset. These results are randomly chosen and not cherry-picked.
Image DINO-B MAE-B MAE-H SD (10 Classes) SD SAM
@@
Figure 9: Qualitative Results on the DRAM (Cohen et al., 2022) dataset. These results are randomly chosen and not cherry-picked.
Image DINO-B MAE-B MAE-H SD (10 Classes) SD SAM
@@
Figure 10: Qualitative Results on the EgoHOS (Zhang et al., 2022b) dataset. These results are randomly chosen and not cherry-picked.
Image DINO-B MAE-B MAE-H SD (10 Classes) SD SAM
@@
Figure 11: Qualitative Results on the iShape (Yang et al., 2021) dataset. These results are randomly chosen and not cherry-picked.
Image DINO-B MAE-B MAE-H SD (10 Classes) SD SAM
@@
Figure 12: Qualitative Results on the PIDRay (Zhang et al., 2022a) dataset. These results are randomly chosen and not cherry-picked.
Refer to caption Refer to caption
DINO-B
MAE-H (5 classes)
MAE-B
MAE-H (10 classes)
MAE-H
MAE-H (COCO)
MAE-H (ClevrTex)
SD (COCO)
SD (ClevrTex)
SD (5 classes)
SD (10 classes)
SD
SAM
Figure 13: Qualitative comparison of all models on a challenging, in-the-wild image.
DINO-B
MAE-B
MAE-H
MAE-H (5 classes)
MAE-H (10 classes)
MAE-H (ClevrTex)
MAE-H (COCO)
SD (5 classes)
SD (10 classes)
SD (ClevrTex)
SD (COCO)
SD
SAM
Figure 14: Qualitative comparison of all models on a challenging, in-the-wild image.
Refer to caption Refer to caption
DINO-B
MAE-H (5 classes)
MAE-B
MAE-H (10 classes)
MAE-H
MAE-H (COCO)
MAE-H (ClevrTex)
SD (COCO)
SD (ClevrTex)
SD (5 classes)
SD (10 classes)
SD
SAM
Figure 15: Qualitative comparison of all models on a challenging, in-the-wild image.
Refer to caption Refer to caption
DINO-B
MAE-H (5 classes)
MAE-B
MAE-H (10 classes)
MAE-H
MAE-H (COCO)
MAE-H (ClevrTex)
SD (COCO)
SD (ClevrTex)
SD (5 classes)
SD (10 classes)
SD
SAM
Figure 16: Qualitative comparison of all models on a challenging, in-the-wild image.
Refer to caption Refer to caption
DINO-B
MAE-H (5 classes)
MAE-B
MAE-H (10 classes)
MAE-H
MAE-H (COCO)
MAE-H (ClevrTex)
SD (COCO)
SD (ClevrTex)
SD (5 classes)
SD (10 classes)
SD
SAM
Figure 17: Qualitative comparison of all models on a challenging, in-the-wild image.
DINO-B
MAE-B
MAE-H
MAE-H (5 classes)
MAE-H (10 classes)
MAE-H (ClevrTex)
MAE-H (COCO)
SD (5 classes)
SD (10 classes)
SD (ClevrTex)
SD (COCO)
SD
SAM
Figure 18: Qualitative comparison of all models on a challenging, in-the-wild image.
Refer to caption Refer to caption
DINO-B
MAE-H (5 classes)
MAE-B
MAE-H (10 classes)
MAE-H
MAE-H (COCO)
MAE-H (ClevrTex)
SD (COCO)
SD (ClevrTex)
SD (5 classes)
SD (10 classes)
SD
SAM
Figure 19: Qualitative comparison of all models on a challenging, in-the-wild image.